OpenStudy (anonymous):

Use Property 2 for radicals to simplify the following radical expression as much as possible.

OpenStudy (anonymous):

\[\sqrt{\frac{ 4 }{ 9 }}\]

OpenStudy (anonymous):

I know, it's 2/3

OpenStudy (anonymous):

@hartnn

OpenStudy (ipwnbunnies):

Err, I assume it's that division one. Well, this term can be rewritten as this: \[\sqrt{\frac{4}{9}} = \frac{\sqrt{4}}{\sqrt{9}}\]

OpenStudy (anonymous):

2/3, right?

OpenStudy (ipwnbunnies):

Yeah, sqrt of 4 is 2. Sqrt of 9 is 3. 2/3. :)

OpenStudy (anonymous):

Thaaank you!!